an item has a regular price of $40 at three different stores. each store has a different offer for the item
1
Expert's answer
2019-10-16T11:14:48-0400

Regular price of item at all the three stores"=40"

Offer at the first store"=10" %

Price at the first store"=40-\frac{10}{100}\times 40=36"

Offer at the second store is double of that in the first store, so

Price at the second store"=40-\frac{2\times 10}{100}\times 40=32"

Offer at the third store is "\frac{3}{4}^{th}" of that in the second store.

Price at the third store"=40-\frac{3 \times 2\times 10}{4\times 100}\times 40=34"

Hence price at the third store is the mean of that in the first and the second store.